OpenStudy (purplemexican):

A ball is dropped from the height of 475 centimeters. Each time it hits the floor, it rebounds to a height that is %24, percent of the previous height. Find the total distance traveled by the ball.

OpenStudy (welshfella):

This can be represented by an infinite geometric series

OpenStudy (purplemexican):

yes can you please help me figure out how to make it a infinite geometric series

OpenStudy (welshfella):

on each bounce it travels same distance up and down so the common ratio will be 2*0.24 = 0.48 First term will be 465 so the series is 465, 465*0.48, 465*0.48^2...

OpenStudy (welshfella):

and the sum to infinity = a1 ------ 1 - r where a1 = first term and r = common ratio.

OpenStudy (purplemexican):

wouldn't the first term be 475

OpenStudy (welshfella):

yes 475 Typo

OpenStudy (purplemexican):

so how would i find the total distance the ball traveled @welshfella

OpenStudy (welshfella):

total distance = sum to infinity ( the formula in my last post)

OpenStudy (welshfella):

just plug in a1 = 475 and r = 0.48

OpenStudy (purplemexican):

\[\frac{ a1 }{ r }\]

OpenStudy (welshfella):

no a1 / (1 - r)

OpenStudy (purplemexican):

sorry i forgot the 1-r

OpenStudy (purplemexican):

so the sum should be 913.461

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

round it to tenths 913.5